OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = –4(x + 6)^2 + 2? (6, 2) (–6, 2) (6, –2) (–6, –2)

OpenStudy (anonymous):

its (-6,2)

hartnn (hartnn):

For y = a(x -h)^2 + k, the vertex is (h,k).

OpenStudy (anonymous):

\[y'=-8(x+6)\]The vertex is the critical point of the parabola.\[0=-8(x+6)\\0=x+6\\x=-6\]Now, just determine the y.\[y=-4(-6+6)^2+2=2\]So the vertex is...\[(-6,2)\]

OpenStudy (anonymous):

Or, the easy way to do it... we know the vertical parabola is in vertex form i.e.\[y-k = a(x-h)^2\]... then the vertex is:\[(h,k)\]In our case, we find the variables are as follows.\[y=4(x+6)^2+2\\y-2=4(x- -6)^2\]\[a=4\\h=-6\\k=2\]So, the vertex is then...\[(h,k)=(-6,2)\]

hartnn (hartnn):

\(\large \frak{\color{red}{WELCOME \quad TO \quad OPENSTUDY, \quad DEANNA !}}\)